OpenStudy (anonymous):

simplify sqrt 72 - 3 sqrt 12 + sqrt 192

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi :)

OpenStudy (anonymous):

I think I figured it out would it be 6 sqrt 2 + 2 sqrt 3

OpenStudy (misty1212):

write each number in simplest radical form first then combine like terms for example \(\sqrt{12}=\sqrt{4\times 3}=\sqrt4\sqrt3=2\sqrt3\) the last line is simplest radical form

OpenStudy (anonymous):

Oh okay! so then it come to 6 sqrt 2 = 2 sqrt 3 right ?

OpenStudy (misty1212):

\[-3\sqrt{12}=-6\sqrt3\]

OpenStudy (anonymous):

that isn't one of the choices though

OpenStudy (misty1212):

that is only one small part that is why

OpenStudy (misty1212):

you have to do the same thing with the other two terms

OpenStudy (anonymous):

Oh okay! so then it would come to 8 sqrt 3 - sqrt 1?

OpenStudy (anonymous):

or no ?